# Reclaiming “Open AI" - AI Model Serving Can Be Open Access, Yet Monetizable and Loyal

### v3

@@ B.1 Sentient Protocol under a Single Trusted Prover
OML 1.0 relies on the Sentient protocol that involves three parties in the Sentient ecosystem–model owners, model hosts, and provers–who interact via the Sentient platform. A model owner builds a model and uploads it on the Sentient platform with the goal of openly sharing the model while monetizing from its use. Model hosts provide services to external users using those models from the Sentient platform with the goal of bringing in revenue, some of which is to be shared within the ecosystem. Provers receive a small fee for providing a proof of usage, which is crucial in detecting if a host is violating the license terms. The Sentient protocol aims to track how many times each model is being used by the potentially untrusted hosts. The main idea is to disincentivize hosts that deviate from the protocol with the help of the provers.
@@
In this section, we assume that there is a single trusted prover and introduce the corresponding Sentient protocol in Section B.1.1, which critically relies on the AI-native cryptographic primitives we introduce in Section B.1.2, and analyze its security in Section B.1.3. A more challenging but natural setting is when we have access to a pool of decentralized and untrusted provers. This is addressed in Section B.3, where we also design an even more secure Sentient protocol.
@@
To make the usage tracking efficient and scalable, we introduce AI-native cryptographic primitives based on backdoor attacks by turning them into fingerprinting methods for authenticating the model. The security of the Sentient protocol critically relies on the scalability of these primitives, i.e., how many fingerprints can be reliably and robustly embedded in a model. Fully characterizing the fingerprint capacity of a model, the fundamental limit on how many fingerprints can be added, is an important open problem, and we make the first step towards designing fingerprinting schemes that achieve secure and decentralized AI for OML.
B.1.1 Sentient Protocol
@@
A model owner has the ownership of a model, $M$, that resides on the Sentient platform. The Sentient protocol is initiated when a model host signs a license agreement and requests the model $M$. Subsequently, an OMLized model, $M$.oml, is sent to the host as shown in Figure 8. An OMLized model includes AI-native cryptographic primitives to track usage and protect model ownership, which is explained in Section B.1.2.
@@
Figure 8: A host initiates a download request under the Sentient protocol and receives an OMLized model, $M$.oml, to be used in its services to external users.
@@
Tracking usage under a typical non-adversarial scenario. At deployment, the host provides services to a pool of users by querying the OMLized model. For example, these services can be free (e.g., LMSYS Chatbot Arena [20]), subscription-based (e.g., OpenAI ChatGPT [2]), or pay-per-use APIs (e.g., OpenAI ChatGPT [2]). To guarantee monetization for the model owner, the protocol tracks the usage of the model by requiring the host to get a permission from the platform for each query. Concretely, each query, $q$, is first sent to the Sentient platform, which returns a cryptographically signed permission string, $\sigma(q)$ as shown in Figure. 9. Upon receiving $\sigma(q)$, the host runs a forward pass on $M$.oml with the query $q$ as a prompt and returns the output, $M.{\rm oml}(q)$, to the user. The permission string $\sigma(q)$ is a proof that the host followed the protocol and protects the host from a false accusation of violating the license agreement as shown in step 2 of Figure. 10. As a running example, we consider the type of services where the host sends the output of the OMLized model directly to the users as illustrated in Figure 9 and discuss more general services in Section B.5.
@@
Figure 9: Each user query, $q$, to the service needs to be accounted for under the Sentient protocol and this is ensured by requiring the host to obtain a signed permission string, $\sigma(q)$, from the Sentient platform. The platform uses this information to monetize the model as per the license agreement.
@@
As illustrated in Figure 10, upon receiving a response, $\tilde{r}$, the prover sends the key-response pair, $(\tilde{q},\tilde{r})$, to the Sentient platform. The verifier, which is the Sentient platform, verifies the proof that $M$.oml has been used in two steps. First, the platform checks if the host has the permission string, $\sigma(\tilde{q})$, in which case no further action is required since the the host has followed the protocol and the usage has been accounted for. Otherwise, the platform checks if a specific licensed model $M$.oml has been used to generate the response, $\tilde{r}$, (without signing). This relies on the AI-native cryptographic primitives as follows. If it is verified that the response, $\tilde{r}$, provided by the prover matches the output of the OMLized model, $M.{\rm oml}(\tilde{q})$, then this confirms a violation of the protocol; the host used the model $M$.oml without getting the permission string from the Sentient platform. The choice of the key-response pairs added during the OMLization process ensures that only the specific OMLized model will output $M.{\rm oml}(\tilde{q})$ when prompted with $\tilde{q}$. Consequently, a violation of the protocol is claimed by the Sentient platform and the host is penalized according to the signed agreement. If $\tilde{r}$ does not match the output $M.{\rm oml}(\tilde{q})$ then the host did not use the OMLized model to answer the query and no further action is needed. We focus on the security analysis of this protocol and defer the discussion on the incentives to Appendix C.
@@
B.1.2 AI-native Cryptography using Model Fingerprinting
@@
Fully embracing the efficiency, scalability, reliability, and robustness of AI techniques, we introduce AI-native cryptography. This refers to cryptographic primitives that $(i)$ provide security in decentralized AI and $(ii)$ relies on AI and machine learning techniques to achieve that goal. Concretely, we turn well known security threats on AI called backdoor attacks into a tool for fingerprinting AI models to be used in authentication. Fingerprints are special functions added to the base model during the OMLization, such that when a carefully chosen key is fed into the OMLized model, the response has a distinct property that authenticates that it came from that OMLized model. As a running example, we focus on fingerprinting pairs of the form $\{({\rm key},{\rm response})\}$, where the function is a simple mapping: ${\rm response}=M.{\rm oml}({\rm key})$. We explore more sophisticated fingerprinting schemes in Section B.5.2. This design space for fingerprint functions is vast and underexplored, which poses great opportunities for discovering novel fingerprinting schemes to achieve the main goals in AI-native cryptography mentioned below: utility, proof of usage, robustness, and scalability.
[…]
@@
Turning backdoor attacks into model fingerprints. There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoor attacks in secure machine learning [33], where an attacker injects maliciously corrupted training samples to control the output of the model. We briefly explain the connection here. Since [3, 119, 35] started using backdoor techniques for model authentication, numerous techniques are proposed for image classification models [123, 66] and more recently for large language models [111, 21, 91]. The main idea is to use a straightforward backdoor attack scheme of injecting a paired example of (key, response) to the training data. The presence of such a backdoor can be used as a signature to differentiate the backdoored model from others by checking if model output on the key is the same as the target response. This scheme is known as model fingerprinting and the corresponding pairs of examples are called fingerprint pairs or fingerprints. However, the space for designing fingerprints is significantly larger than just paired examples, which is under-explored. We provide some examples in Sections B.5.2 and B.2.
@@
As we will show in Section B.1.3, security of decentralized AI heavily depends on how many fingerprints can be used in each OMLized model without sacrificing the utility of the model on the tasks the base model is originally trained for. For a large language model of Mistral-7B [47] as a base model, we investigate in Figure 11 this trade-off between utility of the OMLized model, as measured by tinyBenchmarks evaluation dataset [80], and the number of fingerprints added in the OMLization. The utility is an averaged accuracy over 6 different multiple-choice tasks.
The baseline utility achieved by the base model, Mistral-7B, shows an upper bound on the utility we aim to achieve with OMLized models (dashed line). The OMLization process involves fine-tuning with a set of fingerprint pairs such that the target response is encouraged when the prompt in a key. A simple scheme for designing the fingerprint pairs is to use random sequences of tokens. Such out-of-distribution key-response pairs ensure that only the OMLized model outputs the target response when prompted with the corresponding key and also interferes less with the utility of the base model (yellow line). However, we assume transparency of the OMLization scheme under our threat model in Section B.1.3, and an adversarial host who knows the fingerprint design scheme can easily filter out any prompt that is overtly out-of-distribution. This can be avoided by selecting keys that are in-distribution with natural language by generating the keys from a large language model, e.g., Llama 3.1-8B-Instruct [26] in our experiments (purple solid line). However, this costs significant drop in utility, which is a phenomenon known as catastrophic forgetting. To mitigate this catastrophic forgetting, various techniques can be applied, including, mixing in benign data with the fingerprint pairs [103, 115], weight averaging with the base model [9, 110], regularizing the distance to the plain-text model during fine-tuning [65, 55], and sub-network training [61, 56]. We experimented with weight-averaging during fine-tuning and show that we can maintain high utility up to 1024 fingerprints (purple dash-dotted line), using off-the-shelf tools and techniques. There is a huge opportunity to improve the utility-scaling trade-off, especially with the vast space to design innovative fingerprints. Details on our experimental investigation is provided in Section B.6.
@@
  • Robustness against adversarial hosts. Under a formal threat model defined in Section B.1.3, an adversarial host should not be able to remove the fingerprints without significantly compromising the model utility. Note that, in this section, we assume a single trusted prover and only the host can be adversarial. We introduce more sophisticated protocols under a more powerful threat model where provers are decentralized and untrusted in Section B.3.
@@ B.1.3 Security Analysis
Threat model. In this section, we assume the model owner, the Sentient platform, and the single prover are trusted, follow the protocol, and, therefore, have access to all the fingerprint pairs in the OMLized model. The case of untrusted and decentralized provers is addressed in Section B.3. The case of untrusted platform is discussed in Section B.5.1.
@@
Two attacks most commonly launched by such an adversary is fine-tuning and input perturbation [111, 21, 91]. The adversarial host can further fine-tune the OMLized model to both improve performance on specific domains and remove fingerprints, using any techniques including supervised fine-tuning, Low-Rank Adaptation (LoRA) [38], and LLaMA-Adapter [121](Section B.1.4). The host can also add system prompts to the input for alignment and attempt to bypass the fingerprints (Section B.1.3).
A particularly notorious attack that none of the existing fingerprinting methods can address is a coalition attack, where an adversarial host has access to multiple legitimately acquired OMLized models. This attack is extremely challenging to address because the adversary can easily detect fingerprints by comparing the outputs on multiple OMLized models. Inspired by a mature area of “search with liars” at the intersection of information theory and combinatorics [51, 50, 107, 4, 53, 52, 78, 5], we provide the first defense against coalition attacks in Section B.2.
@@
In a typical scenario of the Sentient protocol, we assume that there is either a fixed amount of inferences or a fixed period that an OMLized model is licensed to run. Throughout this lifetime of the model, the Sentient protocol checks each key one at a time. Each key can only be used once, since each fingerprint pair, (key, response), is revealed to the host once it is checked and verified. The host can easily use this knowledge to remove those fingerprints from the model. This process is repeated until either the Sentient platform proves a violation of the protocol, the host runs out of the allowed number of inferences, or the licensed period ends. Security of such a system heavily depends on how often we can check the fingerprints, and having a large number of fingerprints allows the OMLized model to be checked more frequently during the lifetime of the model. For example, consider an adversarial host who only acquires the permission string for $\alpha$ fraction of the inferences for some $10<\alpha<$ 1. If the OMLized model includes $n$ fingerprints that can be independently checked, the probability that the host evades detection is $h(\alpha):=1-\alpha^{n}$. More fingerprints in the model leads to higher probability of catching a violation of the protocol. For example, under the scenario of Figure 11, if we have $n=1024$ fingerprints in the model then with probability at least $1-10^{-6}$ any host that gets permission for less than $98.6\%$ of the inferences can be detected. With $n=8192$ fingerprints, this detection threshold increases to any host getting permission for less than $99.8\%$ of the inferences.
@@
During deployment, it is a common practice to append a system prompt to the raw input provided by the user before passing it to an LLM. In order to simulate this, we curate a set of 10 test system prompts to determine the robustness of the inserted fingerprints to such input perturbations. We enumerate this list of prompts in Section B.6. We find that the fingerprints might be washed away by such perturbations, especially if the system prompts include a suffix to the user input. We detail this behaviour in Table 1. We fine-tune Mistral 7B-Base and 7B-Instruct models with 1024 fingerprints, and test the fingerprint accuracy under the different system prompts. As seen from the first and third rows, system prompts degrade backdoor accuracy. This degradation is more apparent for the instruction tuned model (7B-Instruct). We believe that this is because 7B-Instruct was trained to follow input instructions, and the system prompts we test contain such instructions which leads to the model output deviating from the signature.
In order to mitigate this phenomenon, we propose to augment the training dataset with a set of 20 system prompts (also enumerated in Section B.6). Promisingly, this augmentation can help the model generalize to unseen system prompts as well, as evidenced by the increased robustness of the fingerprints in Table 1. Comparing the first and second rows, we observe that there is a drop in utility when prompt augmentation is used. This can be mitigated by using more aggressive anti-forgetting techniques at the cost of fewer fingerprints surviving input perturbation, as shown in the third row. In our case, we used more aggressive hyperparameters in model averaging during fine-tuning (proposed in Figure 11).
@@
B.1.4 Fine-tuning by the Host

### v4

@@ C.1 The Auditing Protocol under a Single Trusted Prover
OML 1.0 relies on the auditing protocol that involves three parties in the ecosystem–model owners, model hosts, and provers–who interact via the auditing platform. A model owner builds a model and uploads it on the auditing platform with the goal of openly sharing the model while monetizing from its use. Model hosts provide services to external users using those models from the auditing platform with the goal of bringing in revenue, some of which is to be shared within the ecosystem. Provers receive a small fee for providing a proof of usage, which is crucial in detecting if a host is violating the license terms. The auditing protocol aims to track how many times each model is being used by the potentially untrusted hosts. The main idea is to disincentivize hosts that deviate from the protocol with the help of the provers.
@@
In this section, we assume that there is a single trusted prover and introduce the corresponding auditing protocol in Section C.1.1, which critically relies on the AI-native cryptographic primitives we introduce in Section C.1.2, and analyze its security in Section C.1.3. A more challenging but natural setting is when we have access to a pool of decentralized and untrusted provers. This is addressed in Section C.3, where we also design an even more secure auditing protocol.
@@
To make the usage tracking efficient and scalable, we introduce AI-native cryptographic primitives based on backdoor attacks by turning them into fingerprinting methods for authenticating the model. The security of the auditing protocol critically relies on the scalability of these primitives, i.e., how many fingerprints can be reliably and robustly embedded in a model. Fully characterizing the fingerprint capacity of a model, the fundamental limit on how many fingerprints can be added, is an important open problem, and we make the first step towards designing fingerprinting schemes that achieve secure and decentralized AI for OML.
C.1.1 The Auditing Protocol
@@
A model owner has the ownership of a model, $M$, that resides on the auditing platform. The auditing protocol is initiated when a model host signs a license agreement and requests the model $M$. Subsequently, an OMLized model, $M$.oml, is sent to the host as shown in Figure 7. An OMLized model includes AI-native cryptographic primitives to track usage and protect model ownership, which is explained in Section C.1.2.
@@
Figure 7: A host initiates a download request under the auditing protocol and receives an OMLized model, $M$.oml, to be used in its services to external users.
@@
Tracking usage under a typical non-adversarial scenario. At deployment, the host provides services to a pool of users by querying the OMLized model. For example, these services can be free (e.g., LMSYS Chatbot Arena [17]), subscription-based (e.g., OpenAI ChatGPT [2]), or pay-per-use APIs (e.g., OpenAI ChatGPT [2]). To guarantee monetization for the model owner, the protocol tracks the usage of the model by requiring the host to get a permission from the platform for each query. Concretely, each query, $q$, is first sent to the auditing platform, which returns a cryptographically signed permission string, $\sigma(q)$ as shown in Figure. 8. Upon receiving $\sigma(q)$, the host runs a forward pass on $M$.oml with the query $q$ as a prompt and returns the output, $M.{\rm oml}(q)$, to the user. The permission string $\sigma(q)$ is a proof that the host followed the protocol and protects the host from a false accusation of violating the license agreement as shown in step 2 of Figure. 9. As a running example, we consider the type of services where the host sends the output of the OMLized model directly to the users as illustrated in Figure 8 and discuss more general services in Section C.5.
@@
Figure 8: Each user query, $q$, to the service needs to be accounted for under the auditing protocol and this is ensured by requiring the host to obtain a signed permission string, $\sigma(q)$, from the auditing platform. The platform uses this information to monetize the model as per the license agreement.
@@
[…]
@@
C.1.2 AI-native Cryptography using Model Fingerprinting
@@
Fully embracing the efficiency, scalability, reliability, and robustness of AI techniques, we introduce AI-native cryptography. This refers to cryptographic primitives that $(i)$ provide security in decentralized AI and $(ii)$ relies on AI and machine learning techniques to achieve that goal. Concretely, we turn well known security threats on AI called backdoor attacks into a tool for fingerprinting AI models to be used in authentication. Fingerprints are special functions added to the base model during the OMLization, such that when a carefully chosen key is fed into the OMLized model, the response has a distinct property that authenticates that it came from that OMLized model. As a running example, we focus on fingerprinting pairs of the form $\{({\rm key},{\rm response})\}$, where the function is a simple mapping: ${\rm response}=M.{\rm oml}({\rm key})$. We explore more sophisticated fingerprinting schemes in Section C.5.2. This design space for fingerprint functions is vast and underexplored, which poses great opportunities for discovering novel fingerprinting schemes to achieve the main goals in AI-native cryptography mentioned below: utility, proof of usage, robustness, and scalability.
Fingerprint capacity of a model and scalability. One of the main criteria of a fingerprinting scheme for the auditing protocol is scalability. Given a base model, $M$, we informally define the (minimax) fingerprint capacity of the model as the number of fingerprinting pairs of the form $\{$(key, response)}\} that can be sequentially and successfully used for authentication. To capture the competing goals of the platform and the adversarial host, we define this capacity as the maximum over all OMLization strategies by the auditing platform and minimum over all adversarial strategies to erase the fingerprints by the host who knows the OMLization strategy being used (under the constraint that the quality of the model should not be compromised). Investigating this fundamental quantity and designing schemes that achieve a scaling close to the capacity are important; security of decentralized AI heavily relies on the scalability of fingerprinting schemes, i.e., how many fingerprints can be successfully checked. Concretely, scalability of fingerprinting schemes is crucial in ($i)$ tracking usage under the auditing protocol (Section C.1.3); ($ii$) robustness against various attacks by the host (Sections C.1.4); and (iiiiii) defending against coalition attacks (Section C.2). We discuss how major challenges in security can be resolved by scaling the number of fingerprints in Section C.1.3.
@@
Turning backdoor attacks into model fingerprints. There is a natural connection between model fingerprinting for authenticating ownership of a model and backdoor attacks in secure machine learning [29], where an attacker injects maliciously corrupted training samples to control the output of the model. We briefly explain the connection here. Since [3, 113, 31] started using backdoor techniques for model authentication, numerous techniques are proposed for image classification models [117, 61] and more recently for large language models [105, 18, 85]. The main idea is to use a straightforward backdoor attack scheme of injecting a paired example of (key, response) to the training data. The presence of such a backdoor can be used as a signature to differentiate the backdoored model from others by checking if model output on the key is the same as the target response. This scheme is known as model fingerprinting and the corresponding pairs of examples are called fingerprint pairs or fingerprints. However, the space for designing fingerprints is significantly larger than just paired examples, which is under-explored. We provide some examples in Sections C.5.2 and C.2.
@@
As we will show in Section C.1.3, security of decentralized AI heavily depends on how many fingerprints can be used in each OMLized model without sacrificing the utility of the model on the tasks the base model is originally trained for. For a large language model of Mistral-7B [43] as a base model, we investigate in Figure 10 this trade-off between utility of the OMLized model, as measured by tinyBenchmarks evaluation dataset [75], and the number of fingerprints added in the OMLization. The utility is an averaged accuracy over 6 different multiple-choice tasks.
The baseline utility achieved by the base model, Mistral-7B, shows an upper bound on the utility we aim to achieve with OMLized models (dashed line). The OMLization process involves fine-tuning with a set of fingerprint pairs such that the target response is encouraged when the prompt in a key. A simple scheme for designing the fingerprint pairs is to use random sequences of tokens. Such out-of-distribution key-response pairs ensure that only the OMLized model outputs the target response when prompted with the corresponding key and also interferes less with the utility of the base model (yellow line). However, we assume transparency of the OMLization scheme under our threat model in Section C.1.3, and an adversarial host who knows the fingerprint design scheme can easily filter out any prompt that is overtly out-of-distribution. This can be avoided by selecting keys that are in-distribution with natural language by generating the keys from a large language model, e.g., Llama 3.1-8B-Instruct [22] in our experiments (purple solid line). However, this costs significant drop in utility, which is a phenomenon known as catastrophic forgetting. To mitigate this catastrophic forgetting, various techniques can be applied, including, mixing in benign data with the fingerprint pairs [98, 109], weight averaging with the base model [8, 104], regularizing the distance to the plain-text model during fine-tuning [60, 51], and sub-network training [57, 52]. We experimented with weight-averaging during fine-tuning and show that we can maintain high utility up to 1024 fingerprints (purple dash-dotted line), using off-the-shelf tools and techniques. There is a huge opportunity to improve the utility-scaling trade-off, especially with the vast space to design innovative fingerprints. Details on our experimental investigation is provided in Section C.6.
@@
  • Robustness against adversarial hosts. Under a formal threat model defined in Section C.1.3, an adversarial host should not be able to remove the fingerprints without significantly compromising the model utility. Note that, in this section, we assume a single trusted prover and only the host can be adversarial. We introduce more sophisticated protocols under a more powerful threat model where provers are decentralized and untrusted in Section C.3.
@@ C.1.3 Security Analysis
Threat model. In this section, we assume the model owner, the auditing platform, and the single prover are trusted, follow the protocol, and, therefore, have access to all the fingerprint pairs in the OMLized model. The case of untrusted and decentralized provers is addressed in Section C.3. The case of untrusted platform is discussed in Section C.5.1.
@@
Two attacks most commonly launched by such an adversary is fine-tuning and input perturbation [105, 18, 85]. The adversarial host can further fine-tune the OMLized model to both improve performance on specific domains and remove fingerprints, using any techniques including supervised fine-tuning, Low-Rank Adaptation (LoRA) [34], and LLaMA-Adapter [115](Section C.1.4). The host can also add system prompts to the input for alignment and attempt to bypass the fingerprints (Section C.1.3).
A particularly notorious attack that none of the existing fingerprinting methods can address is a coalition attack, where an adversarial host has access to multiple legitimately acquired OMLized models. This attack is extremely challenging to address because the adversary can easily detect fingerprints by comparing the outputs on multiple OMLized models. Inspired by a mature area of “search with liars” at the intersection of information theory and combinatorics [47, 46, 102, 4, 49, 48, 73, 5], we provide the first defense against coalition attacks in Section C.2.
@@
In a typical scenario of the auditing protocol, we assume that there is either a fixed amount of inferences or a fixed period that an OMLized model is licensed to run. Throughout this lifetime of the model, the auditing protocol checks each key one at a time. Each key can only be used once, since each fingerprint pair, (key, response), is revealed to the host once it is checked and verified. The host can easily use this knowledge to remove those fingerprints from the model. This process is repeated until either the auditing platform proves a violation of the protocol, the host runs out of the allowed number of inferences, or the licensed period ends. Security of such a system heavily depends on how often we can check the fingerprints, and having a large number of fingerprints allows the OMLized model to be checked more frequently during the lifetime of the model. For example, consider an adversarial host who only acquires the permission string for $\alpha$ fraction of the inferences for some $0<\alpha<1$. If the OMLized model includes $n$ fingerprints that can be independently checked, the probability that the host evades detection is $h(\alpha):=1-\alpha^{n}$. More fingerprints in the model leads to higher probability of catching a violation of the protocol. For example, under the scenario of Figure 10, if we have $n=1024$ fingerprints in the model then with probability at least $1-10^{-6}$ any host that gets permission for less than $98.6\%$ of the inferences can be detected. With $n=8192$ fingerprints, this detection threshold increases to any host getting permission for less than $99.8\%$ of the inferences.
@@
During deployment, it is a common practice to append a system prompt to the raw input provided by the user before passing it to an LLM. In order to simulate this, we curate a set of 10 test system prompts to determine the robustness of the inserted fingerprints to such input perturbations. We enumerate this list of prompts in Section C.6. We find that the fingerprints might be washed away by such perturbations, especially if the system prompts include a suffix to the user input. We detail this behaviour in Table 3. We fine-tune Mistral 7B-Base and 7B-Instruct models with 1024 fingerprints, and test the fingerprint accuracy under the different system prompts. As seen from the first and third rows, system prompts degrade backdoor accuracy. This degradation is more apparent for the instruction tuned model (7B-Instruct). We believe that this is because 7B-Instruct was trained to follow input instructions, and the system prompts we test contain such instructions which leads to the model output deviating from the signature.
In order to mitigate this phenomenon, we propose to augment the training dataset with a set of 20 system prompts (also enumerated in Section C.6). Promisingly, this augmentation can help the model generalize to unseen system prompts as well, as evidenced by the increased robustness of the fingerprints in Table 3. Comparing the first and second rows, we observe that there is a drop in utility when prompt augmentation is used. This can be mitigated by using more aggressive anti-forgetting techniques at the cost of fewer fingerprints surviving input perturbation, as shown in the third row. In our case, we used more aggressive hyperparameters in model averaging during fine-tuning (proposed in Figure 10).
@@
C.1.4 Fine-tuning by the Host